Show that u(x,y)=excosy and v(x,y)=exsiny are harmonic for all values of (x,y).


1
Expert's answer
2021-09-28T10:37:00-0400

Let us show that "u(x,y)=e^x\cos y" and "v(x,y)=e^x\sin y" are harmonic for all values of "(x,y)." Taking into account that "\frac{\partial u(x,y)}{\partial x}=e^x\cos y,
\frac{\partial u^2(x,y)}{\partial x^2}=e^x\cos y,
\frac{\partial u(x,y)}{\partial y}=-e^x\sin y,
\frac{\partial u^2(x,y)}{\partial y^2}=-e^x\cos y,"

we conclude that "\frac{\partial u^2(x,y)}{\partial x^2}+\frac{\partial u^2(x,y)}{\partial y^2}=e^x\cos y-e^x\cos y=0," and hence the function "u(x,y)=e^x\cos y" is harmonic.


Since "\frac{\partial v(x,y)}{\partial x}=e^x\sin y,
\frac{\partial v^2(x,y)}{\partial x^2}=e^x\sin y,
\frac{\partial v(x,y)}{\partial y}=e^x\cos y,
\frac{\partial v^2(x,y)}{\partial y^2}=-e^x\sin y,"

we conclude that "\frac{\partial v^2(x,y)}{\partial x^2}+\frac{\partial v^2(x,y)}{\partial y^2}=e^x\sin y-e^x\sin y=0," and hence the function "v(x,y)=e^x\sin y" is harmonic.